OpenStudy (anonymous):

The product of two consecutive even integers is 10 less than 5 times their sum. Find the two integers.

OpenStudy (anonymous):

write the equation and then solve it. i did it in my head so it's not too hard... the question, that is. my head is pretty hard.

OpenStudy (anonymous):

lol How do I write the equation?

OpenStudy (anonymous):

8+10=18, 18*5= 90 8*10=80

OpenStudy (anonymous):

let x and x+2 be the two consecutive integers. then x(x+2)+10=5(x+(x+2))

OpenStudy (anonymous):

That looks sooo confusing lol

OpenStudy (anonymous):

\(x^2+2x+10=10x+10\) solve for x. you should get two answers...

OpenStudy (anonymous):

8 & 10

OpenStudy (anonymous):

nope, that's only one set of answers... there's another

OpenStudy (anonymous):

My online work says I need 2 different sets of numbers, lower & higher values

OpenStudy (anonymous):

ooooh, uhh...I don't know lol

OpenStudy (anonymous):

Do I replace 8 & 10 with x or no?

OpenStudy (anonymous):

well, learn. it's pretty easy if you just give it a shot.

OpenStudy (anonymous):

get the x's on one side of the equation. it simplifies and then factor. set each factor equal to 0 and solve for x. you can do it!

OpenStudy (anonymous):

Do I combine like terms? like 2x & 10x & both 10s?

OpenStudy (anonymous):

subtract 10x and 10 from both sides... what do you get?

OpenStudy (anonymous):

x^2-8x?

OpenStudy (anonymous):

=?

OpenStudy (anonymous):

x^2-8x=0

OpenStudy (anonymous):

yep, now factor the left hand side

OpenStudy (anonymous):

4 & 2? I don't know which is negative.

OpenStudy (anonymous):

\(x^2-8x=0\) what is common to both terms on the left?

OpenStudy (anonymous):

x x(x-8)?

OpenStudy (anonymous):

there you go. now set each factor equal to 0 and solve for x

OpenStudy (anonymous):

x=0 & x-8=0?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so, it's 0 & 8?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

my online work says 8 is wrong :/

OpenStudy (anonymous):

remember. x and x+2 are the two consecutive even integers. as orangemaster pointed out, 8 and 10 are one set of consecutive even integers 0 and 2 are the others

OpenStudy (anonymous):

ooooooh, that's why lol

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thank you! Can you help me with the next one too?